# LogitSpec: Accelerating Retrieval-based Speculative Decoding via Next Next Token Speculation

### v1

@@ Appendix C Pseudo Code to Prepare Attention Inputs
We give a pseudo code to organize the retrieval multiple draft token sequences into a draft tree and prepare its attention mask. More detailed implementations can be found in our attached code.
@@ Appendix D More Implementation Details
D.1 Retrieval Process
The retrieval process of retrieval-based SD methods will affect the overall acceleration. While the existing method PLD retrievals reference with string matching, it is unaffordable for LogitSpec to retrieval in this way. Specifically, suppose the matching ngram size is $m$, the length of the whole token sequence is $n$, string matching takes the time complexity of $O(mn)$. However, as LogitSpec needs to retrieval for all the next next tokens, the retrieval takes the time complexity of $O(kmn)$, where $k$ is the number of the next next tokens.
To address this issue, we propose a simple yet effective method to reduce the retrieval overhead by constructing a hash table. Specifically, we go through the token sequence, storing each key ngram with size $1$ to $m$ and its value ngram (its following tokens). In this way, the retrieval overhead is reduced to $O(n+k)$ with an additional memory cost $O(mn)$. As the length of the token sequence is relatively small, the additional memory cost is negligible. Furthermore, we will update the retrieval model $\mathcal{R}$ with the decoded tokens in the same way during each decoding step.
D.2 Evaluation Instructions
@@
For CNN/DM, we add a simple instruction “Summarize: ” to the instruction:
@@
For GSM8K, we follow the setting with [Liu et al., 2025] and use a 8-shot CoT for inference:
@@ D.3 Dataset Configurations
In our experiments, we evaluate the effectiveness of our LogitSpec on 4 categories of text generation tasks, including Spec-Bench, CNN/DM, GSM8K, and HumanEval. For Spec-Bench and HumanEval, we use the full data for evaluation. For CNN/DM and GSM8K, we randomly sample 1000 questions for evaluation. The maximal generation length is set as 1024 across all the experiments.
D.4 Model Configurations
@@
In our experiments, all the models are deployed with precision float16. The <|eos token|> is set the same as the tokenizer’s <|eos token|>. To effectively alleviate the overhead of rollbacking KV cache to the accepted draft tokens, we follow Medusa [Cai et al., 2024] to allocate continuous GPU memory for all the KV cache.
Appendix E More Experiment Results
Here, we present the results of Llama 2 chat series models on Spec-Bench, CNN/DM, GSM8K, and HumanEval in Table 5 and Table 6.
@@ Appendix E More Experiment Results
Table 5: Experimental results of LogitSpec on CNN/DM [Nallapati et al., 2016], GSM8K [Cobbe et al., 2021] and HumanEval [Chen et al., 2021] with Llama-2. We report the mean accepted tokens per decoding step (MAT) and overall speedup ratio. We bold the best results and underline the suboptimal results for each backbone model.
@@
Table 6: Experimental results of LogitSpec on Spec-Bench [Xia et al., 2024] with Llama-2. We report the speedup ratio on each sub task, mean accepted tokens per decoding step (MAT) and overall speedup ratio. We bold the best results and underline the suboptimal results for each backbone model.
@@
Appendix F Broader Impact
Social Impact. This paper investigates a retrieval-based speculative decoding technique for LLM inference acceleration, enabling more discussions and potential efficient discovery of novel methods. Our LogitSpec enhances LLMs’ efficiency on downstream tasks and potentially improves user’s satisfaction on LLM-based applications.

### v2

@@ Appendix C Pseudo Code to Prepare Attention Inputs
We provide pseudo code to organize the retrieved multiple draft token sequences into a draft tree and prepare its attention mask. More detailed implementations can be found in our attached code.
@@ Appendix D More Implementation Details
D.1 Baselines
We compare LogitSpec against strong speculative decoding baselines. For draft-model-free, plug-and-play methods, we include: (a) Lookahead Decoding (Fu et al., 2024), which utilizes dropped draft tokens as a retrieval corpus; (b) REST (He et al., 2024), which employs an external knowledge base as the retrieval model; and (c) PLD (Saxena, 2023), which extracts n-grams from the current user-input prompt and decoded tokens as the retrieval model. In addition, we report results for a draft-model-based baseline, vanilla speculative sampling (SpS) (Leviathan et al., 2023; Chen et al., 2023) with a small Vicuna-68M draft model, to illustrate the trade-off between extra model parameters / training and acceleration. We use Vicuna series models (7B, 13B, 33B), Llama-2 series models (7B, 13B, 70B), Llama-3.1-Instruct-8B and Qwen-3-8B as backbones, and report results with baseline methods on Vicuna and Llama-2 as the main results, where the performance of baseline methods can be reproduced with their official implementations.
D.2 Hardware and Software Configurations
We follow the setting of Spec-Bench with Pytorch (Paszke et al., 2019) 2.6.0 and CUDA (NVIDIA et al., 2020) 12.4. Our implementation is based on Hugging Face transformers. For main experiments with Vicuna and Llama-2 models (Section 5.1 and all baseline comparisons), we use transformers (Wolf et al., 2020) version 4.37.1. For additional experiments with Llama-3.1-8B-Instruct and Qwen3-8B (Section E.1), we use transformers version 4.57.1, which is necessary to support these newer model architectures. Experiments with models between 7B and 33B are conducted on a single NVIDIA SXM A100 80G GPU, while experiments with 70B models are conducted on 2 NVIDIA SXM A100 80G GPUs. The inference precision is float16. Following prior work, we conduct experiments with temperature 0 and batch size 1. The maximal generation lengths are 1024.
D.3 Retrieval Process
The retrieval process significantly affects overall acceleration. A naive implementation would be computationally prohibitive for LogitSpec.
Naive Approach.
For each of $k$ speculated next-next tokens, scan the entire prompt (length $L$) to find matching $m$-grams, yielding $O(k\cdot m\cdot L)$ complexity.
Our Hash-Based Optimization.
We reduce this to $O(n+k)$ per decoding step:
  1. 1.
    Preprocessing (amortized): Slide a window of size $m$ over the prompt once, building a hash table mapping each $m$-gram (key) to its continuation span (up to $n$ tokens as value). Cost: $O(L)$, done once per prompt.
  2. 2.
    Per-step retrieval: For each of $k$ next-next token candidates:
    • Construct the $m$-gram query: $O(m)\approx O(1)$ (since $m$ is small, typically 3)
    • Hash table lookup: $O(1)$
    • Copy continuation span: $O(n)$
    Total per step: $O(k\cdot(1+n))=O(k+kn)$.
Memory Cost.
The hash table requires $O(mL)$ memory, which is negligible compared to KV cache (typically several GB for long contexts).
Empirical Validation.
Figure 5 shows retrieval accounts for only 1.17% of total wall-clock time, confirming our optimization is effective in practice.
D.4 Evaluation Instructions
@@
For CNN/DM, we prepend “Summarize: ” to the question:
@@
For GSM8K, we follow the setting with (Liu et al., 2025) and use an 8-shot CoT for inference:
@@ D.5 Dataset Configurations
In our experiments, we evaluate LogitSpec on four categories of text generation tasks, including Spec-Bench, CNN/DM, GSM8K, and HumanEval. For Spec-Bench and HumanEval, we use the full data for evaluation. For CNN/DM and GSM8K, we randomly sample 1000 questions for evaluation. The maximal generation length is set as 1024 across all the experiments.
D.6 Model Configurations
@@
In our experiments, all the models are deployed with precision float16. The <|eos token|> matches the tokenizer’s <|eos token|>. To effectively alleviate the overhead of rolling back the KV cache to the accepted draft tokens, we follow Medusa (Cai et al., 2024) to allocate continuous GPU memory for all the KV cache.
Ethics Statement
This work adheres to the ACL Code of Ethics. Our study does not involve human subjects or personally identifiable information, and we only use publicly available datasets under their respective licenses. We transparently report our methods and potential risks and do not recommend deployment in high-stakes settings without further safety assessments.
Appendix E More In-depth Analysis of LogitSpec
E.1 More Results on Different LLM Backbones
Comparison with State-of-the-Art Retrieval-Based Methods.
To position LogitSpec against recent advances in retrieval-based speculative decoding, we compare with SAM Decoding (Hu et al., 2024) and Token Recycling (Luo et al., 2024). Both methods are plug-and-play and do not require draft models, making them directly comparable to LogitSpec.
[…]
  • vs. Token Recycling: Token Recycling reuses previously rejected draft tokens via an adjacency matrix, which is query-dependent and requires cross-query information. In contrast, LogitSpec is query-independent and uses the target model’s own predictive capability to guide retrieval, making it more robust when query patterns are diverse.
  • vs. SAM Decoding: SAM uses suffix automata for efficient n-gram matching from a static corpus. While this improves matching speed, it still faces the fundamental challenge of low match rates when context repetition is limited. LogitSpec addresses this by speculating next-next tokens, effectively expanding the retrieval space beyond observed n-grams.
[…]
Here, we provide additional results for different LLM backbones to provide more insights into LogitSpec, including Llama 2 chat series models, LLaMA-3.1-8B-Instruct, and Qwen3-8B.
We present the results of Llama 2 chat series models on Spec-Bench, CNN/DM, GSM8K, and HumanEval in Table 7 and Table 8. LogitSpec consistently achieves the best acceleration among all baselines across Spec-Bench, CNN/DM, GSM8K, and HumanEval, further validating its effectiveness. For completeness, Table 5 reports the per-subtask Spec-Bench breakdown on Vicuna 7B/13B/33B, complementing the summary statistics in the main text.
E.2 Per-Task Spec-Bench Breakdown on Vicuna
Per-task Analysis.
The detailed breakdown reveals important insights into LogitSpec’s robustness across diverse tasks. In the Translation task, where almost no reference is available to serve as draft tokens, PLD and Lookahead achieve poor acceleration. Even equipped with a large external database ($\approx$ 12 GB), REST only achieves $1.17\times\sim 1.31\times$ speedup, while LogitSpec achieves $1.38\times\sim 1.43\times$ speedup without external databases, demonstrating the prediction ability of the last logit on the next next token. In contrast, on tasks with high context repetition such as Summarization and RAG, LogitSpec achieves even stronger performance, highlighting how next-next token speculation complements retrieval-based approaches across varying task characteristics.
Models Method MT Trans Sum QA Math RAG MAT Speedup
Vicuna 7B Lookahead 1.40 1.14 1.19 1.24 1.55 1.09 1.66 1.27
REST 1.63 1.31 1.36 1.66 1.21 1.73 1.82 1.48
PLD 1.64 1.04 2.43 1.14 1.61 1.71 1.73 1.59
SpS 1.66 1.13 1.62 1.49 1.47 1.55 2.28 1.49
LogitSpec 1.92 1.39 2.68 1.70 2.22 1.87 2.44 2.01
Vicuna 13B Lookahead 1.30 1.06 1.20 1.12 1.48 1.12 1.63 1.22
REST 1.52 1.17 1.37 1.53 1.19 1.55 1.82 1.38
PLD 1.47 1.02 2.19 1.03 1.57 1.71 1.68 1.48
SpS 1.60 1.13 1.68 1.39 1.53 1.67 2.18 1.49
LogitSpec 1.89 1.43 2.33 1.43 2.23 1.93 2.32 1.93
Vicuna 33B Lookahead 1.32 1.08 1.20 1.06 1.54 1.15 1.61 1.24
REST 1.63 1.27 1.45 1.61 1.30 1.61 1.80 1.48
PLD 1.44 1.06 2.00 1.07 1.55 1.45 1.55 1.42
SpS 1.75 1.28 1.76 1.53 1.69 1.68 2.01 1.61
LogitSpec 1.77 1.38 2.15 1.37 2.00 1.69 2.13 1.76
Table 6 compares LogitSpec against SAM Decoding and Token Recycling on AIME, LongBench, and MATH with the Llama-3.1-8B-Instruct backbone under identical decoding settings ($T=0$, $K=64$, greedy sampling). LogitSpec consistently attains the best MAT and speedup across all datasets.
Table 6: Comparison with stronger retrieval-based baselines on Llama-3.1-8B-Instruct. We report MAT and end-to-end speedup on AIME, LongBench, and MATH, as well as the macro average. All methods share the same hardware, decoding parameters (temperature $T=0$), and tree capacity $K=64$.
Method AIME LongBench MATH Average
MAT Speedup MAT Speedup MAT Speedup MAT Speedup
[…]
We also present more results applying LLaMA-3.1-8B-Instruct and Qwen3-8B on Spec-Bench, CNN/DM, GSM8K, and HumanEval in Tables 9 and 10. We observe that compared to standard autoregressive decoding, LogitSpec consistently achieves around $2\times$ inference acceleration on average across these datasets.
@@ E.2 Per-Task Spec-Bench Breakdown on Vicuna
Table 7: Experimental results of LogitSpec on CNN/DM (Nallapati et al., 2016), GSM8K (Cobbe et al., 2021) and HumanEval (Chen et al., 2021) with Llama-2. We report the mean accepted tokens per decoding step (MAT) and overall speedup ratio. We bold the best results and underline the suboptimal results for each backbone model.
@@
Table 8: Experimental results of LogitSpec on Spec-Bench (Xia et al., 2024) with Llama-2. We report the speedup ratio on each sub task, mean accepted tokens per decoding step (MAT) and overall speedup ratio. We bold the best results and underline the suboptimal results for each backbone model.
@@
Table 9: Experimental results of LogitSpec on Spec-Bench (Xia et al., 2024) with LLaMA-3.1-8B-Instruct and Qwen3-8B. We report the speedup ratio on each sub task, mean accepted tokens per decoding step (MAT) and overall speedup ratio. We bold the best results for each backbone model.
Model Method MT Trans Sum QA MATH RAG MAT Speedup
Llama-3.1-8B-Instruct Vanilla 1.00 1.00 1.00 1.00 1.00 1.00 1.00 1.00
Llama-3.1-8B-Instruct LogitSpec 1.89 1.67 1.94 1.68 2.01 1.77 2.11 1.88
Qwen-3-8B Vanilla 1.00 1.00 1.00 1.00 1.00 1.00 1.00 1.00
Qwen-3-8B LogitSpec 1.71 1.74 1.64 1.65 1.89 1.68 1.95 1.75
Table 10: Experimental results of LogitSpec on CNN/DM (Nallapati et al., 2016), GSM8K (Cobbe et al., 2021) and HumanEval (Chen et al., 2021) with LLaMA-3.1-8B-Instruct and Qwen3-8B. We report the mean accepted tokens per decoding step (MAT) and overall speedup ratio. We bold the best results for each backbone model.
Model Method CNN/DM GSM8K Humaneval Overall
MAT Speedup MAT Speedup MAT Speedup MAT Speedup
Llama-3.1-8B-Instruct Vanilla 1.00 1.00 1.00 1.00 1.00 1.00 1.00 1.00
Llama-3.1-8B-Instruct LogitSpec 2.04 1.85 2.18 1.95 2.63 2.31 2.28 2.04
Qwen-3-8B Vanilla 1.00 1.00 1.00 1.00 1.00 1.00 1.00 1.00
Qwen-3-8B LogitSpec 1.77 1.59 2.18 1.94 2.18 1.92 2.04 1.82
E.3 More Results on Different Benchmarks
Table 11: Results of LogitSpec on AIME datasets.
Method AIME24 AIME25
MAT Speedup MAT Speedup
Vanilla 1.00 1.00 1.00 1.00
LogitSpec 3.41 3.25 3.76 3.33
We provide more results for LogitSpec on complex math reasoning tasks including MATH (Hendrycks et al., 2021) and AIME 24 & 25 (Mathematical Association of America, 2025) datasets in Figure 6 and Table 11. For each of these benchmarks, we report the real-world speedup for each subset, the overall mean accepted tokens per decoding step (MAT), and the overall speedup. These expanded experimental results further corroborate LogitSpec’s effectiveness and applicability, yielding an overall MAT of 3.32 and an overall speedup of $2.78\times$ on MATH, and up to 3.76 MAT and $3.33\times$ speedup on AIME25. This demonstrates LogitSpec’s robust performance even on challenging reasoning tasks.
Figure 6: LogitSpec performance on MATH benchmark subsets.
E.4 Sampling with Non-Zero Temperatures
Why LogitSpec Remains Effective Under Sampling.
A key question is whether LogitSpec’s reliance on top-$k$ logits makes it sensitive to temperature scaling. We clarify the mechanism and provide empirical validation.
Candidate Selection is Temperature-Invariant. In our implementation, speculative next-next token candidates are selected from the raw logits before applying temperature. Since temperature scaling is a monotonic transformation, it preserves ranking:
$z_{i}>z_{j}\quad$ $\Longleftrightarrow\quad\exp(z_{i}/T)>\exp(z_{j}/T)$
$\Longleftrightarrow\quad p_{i}(T)>p_{j}(T),\quad\forall T>0$ (9)
where $p(T)=\text{softmax}(z/T)$. Therefore, the top-$k$ candidate set used for speculation is identical across all temperatures and does not “fill with noise tokens” as $T$ increases.
What Changes: Acceptance Probability. Higher temperatures make the actual sampled trajectory deviate more from the most-likely path, reducing acceptance rates. This is an inherent property of all speculative decoding methods under sampling, not specific to LogitSpec.
To study the robustness of LogitSpec under stochastic decoding, we evaluate performance across temperatures $T\in\{0.2,0.4,0.6,0.8\}$ on AIME, LongBench, and MATH with Llama-3.1-8B-Instruct. As shown in Table 12, LogitSpec maintains strong performance even at higher temperatures, still achieving speedups above $1.8\times$ even at $T=0.8$. This demonstrates that LogitSpec remains effective across diverse generation scenarios.
Table 12: LogitSpec under different sampling temperatures on Llama-3.1-8B-Instruct. Draft candidates are drawn from raw logits before temperature scaling, so the speculative candidate set is unchanged across $T$; higher $T$ only changes the sampled path.
Temperature AIME MAT AIME Speedup LongBench MAT LongBench Speedup MATH MAT MATH Speedup
[…]
E.5 Cross-Lingual Next-Next Token Coverage
To validate the universality of the last logit’s predictive capability across different linguistic structures, we conduct experiments on the MMMLU benchmark covering five languages: English, Chinese, Japanese, German, and French. Table 13 reports the fraction of ground-truth next-next tokens appearing in the top-$k$ logits for each language. The results show consistent coverage across all languages: at $k=64$, over $50\%$ of next-next tokens fall within the top-$k$ predictions for every language. This demonstrates that our observation about the last logit’s predictive capability is not restricted to English or specific linguistic structures, but rather represents a fundamental property of modern LLMs that generalizes across diverse languages with different word orders and morphological characteristics.
Table 13: Fraction of ground-truth next-next tokens that appear within the top-$k$ logits of the last token on MMMLU subsets (Llama-3.1-8B-Instruct). Each language averages 20 questions.
Language $k=4$ $k=8$ $k=16$ $k=32$ $k=64$
English 15.25% 22.02% 33.60% 42.06% 51.83%
Chinese 17.14% 24.21% 36.29% 45.68% 55.07%
Japanese 15.16% 24.30% 34.72% 46.17% 55.87%
German 14.64% 21.44% 32.87% 42.42% 50.64%
French 14.86% 21.77% 32.21% 42.78% 52.22%
These results validate a key claim from our rebuttal: the last logit’s predictive capability is not restricted to English or specific linguistic structures. The consistency across languages with markedly different characteristics—Chinese (logographic, subject-verb-object), Japanese (agglutinative, subject-object-verb), German (inflectional, flexible word order), and French (Romance, subject-verb-object)—suggests this is a fundamental property of modern LLMs’ internal representations rather than a language-specific artifact. Combined with our diverse task coverage (reasoning, QA, summarization, translation, code generation), this cross-lingual validation provides strong evidence for the generality of our core observation.
E.6 More Results on Long-Context Scenarios
[…]
Figure 7: LogitSpec performance on LongBench subsets.
E.7 Ablation Study on Different Pruning Strategies
In Section 4.2, we apply a heuristic pruning strategy to control the number of draft tokens. The role of the pruning algorithm is to effectively control the size of the resulting draft tree, keeping the decoding overhead low without sacrificing too many possibilities, which is a strategy to balance breadth and cost. To provide more insight into the pruning strategy, we conduct ablation studies in Table 14. Specifically, we consider two different pruning strategies:
  • (a)
    Strategy $1$: A rank-based heuristic: if the rank is <<44, we preserve $5$ tokens; <<88, 44 tokens; <<$16$, 33 tokens; <<32, 22 tokens; else $1$ token.
  • (b)
    Strategy $2$: A simple heuristic: preserving $4$ tokens for all speculated next next tokens.
The results in Table 14 demonstrate that while different pruning strategies have some effect on the final performance, the overall performance of LogitSpec is quite robust. The speedup remains stable at approximately $1.9\times$ across these different approaches.
We also would like to clarify that the core contribution of our work is next next token speculation using the last logit to guide and enhance the efficiency and accuracy of retrieval-based speculative decoding. The primary purpose of the pruning algorithm is to serve as an auxiliary module for our core mechanism.
Table 14: Ablation study on pruning strategies. LogitSpec $_{with\hskip 2.84526pts1}$ and LogitSpec $_{with\hskip 2.84526pts2}$ denote LogitSpec with Strategy 1 and Strategy 2, respectively.
Method MT Trans Sum QA MATH RAG MAT Speedup
LogitSpec $_{with\hskip 2.84526pts1}$ 1.90 1.68 1.94 1.66 2.04 1.78 2.12 1.88
LogitSpec $_{with\hskip 2.84526pts2}$ 1.86 1.67 1.87 1.68 2.03 1.74 2.02 1.85
LogitSpec (ori) 1.89 1.67 1.94 1.68 2.01 1.77 2.11 1.88
E.8 Ablation Study on Draft Tree Capacity
As mentioned in Section 4.2, we set the capacity of the draft tree to K=64. To provide further insight, we evaluate a range of $K$ values from $32$ to $128$ in Table 15. These results reveal a clear trade-off: as K increases, the MAT per step improves (from $2.03$ to $2.30$), since a larger tree offers more opportunities for token acceptance. However, the overall speedup ratio peaks at K=64 ($1.92\times$) and subsequently declines. This is because verifying an overly large draft tree introduces significant computational overhead that ultimately negates the advantages of a higher acceptance rate.
Table 15: Ablation study on pruning hyperparameters $K$ using LLaMA-3.1-8B-Instruct as the backbone model.
Method MT Trans Sum QA MATH RAG MAT Speedup
Vanilla 1.00 1.00 1.00 1.00 1.00 1.00 1.00 1.00
LogitSpec 32 1.79 1.62 1.82 1.56 1.93 1.67 1.95 1.79
LogitSpec 48 1.86 1.66 1.90 1.67 2.00 1.73 2.04 1.85
LogitSpec 64 1.89 1.67 1.94 1.68 2.01 1.77 2.11 1.88
LogitSpec 80 1.85 1.64 1.86 1.67 1.84 1.74 2.14 1.83
LogitSpec 96 1.81 1.65 1.83 1.66 1.84 1.73 2.17 1.80
LogitSpec 112 1.79 1.61 1.77 1.64 1.83 1.69 2.21 1.77
LogitSpec 128 1.74 1.62 1.73 1.58 1.81 1.62 2.22 1.73
E.9 Real-World Example
To provide more insights into LogitSpec, we present a real-world example to illustrate how “next next token speculation” allows LogitSpec to succeed where standard retrieval methods fail.
Taking the prefix
Q: A pen costs as much as a pencil and eraser combined. A pencil costs \$1.20 and an eraser costs \$0.30. How much will 8 pens cost?
A: To find the cost of 8 pens, we first need to find the cost of one pen. A pencil costs \$1.20 and an eraser costs \$0.30. The
At the first step, LogitSpec generates the next token “combined” and then speculates the next next token with the following high-probability candidates:
  • total
  • cost
  • combination
  • pen
  • eraser
Then, LogitSpec extends each candidate with retrieved n-grams and verifies these draft sequences:
  • $\times$] . A pencil costs
  • $\times$] total they had
  • $\checkmark$] cost of one pen
  • $\times$] combination (no matched n-grams retrieved)
  • $\times$] pen.
  • $\times$] eraser costs \$
Finally, with the guidance of the last logit, we successfully accept $3$ draft tokens and generate “the combined cost of one pen”. In the next decoding step, LogitSpec again accepts $3$ draft tokens and generates “the combined cost of one pencil and one eraser”. However, without last-logit speculation, standard retrieval can only retrieve the first n-gram “A pencil costs”.
Appendix F Discussion: Integration with Production Frameworks
vLLM Integration Feasibility.
While our current implementation uses Hugging Face transformers, several reviewers asked about integration with production inference frameworks like vLLM. We discuss technical feasibility:
Tree Attention Support. vLLM has recently added native tree attention support:
Compatibility with vLLM’s Architecture. LogitSpec’s draft tree can be naturally represented as multiple sequences with shared prefixes, aligning with vLLM’s batched inference paradigm. Each branch corresponds to a different draft sequence, similar to the prefill process for multiple sequences. The nano-PEARL repository demonstrates a similar integration approach for tree-based verification in vLLM.
Minimal Overhead. Beyond tree attention, LogitSpec only requires:
  • The last logit at each step (already computed by vLLM)
  • CPU-based retrieval and tree construction (can run in parallel with GPU inference)
We conclude that there are no fundamental algorithmic obstacles to deploying LogitSpec in vLLM; the remaining work is engineering integration, which we leave to future work.
Appendix G LLM Usage
We used a large language model (LLM)–based writing assistant solely for grammar and wording improvements on draft text. The LLM did not generate research ideas, claims, proofs, algorithms, code, figures, or analyses, and it did not have access to any non-public data. All edits suggested by the LLM were manually reviewed and either accepted or rewritten by the authors, who take full responsibility for the final content. The LLM is not an author of this paper.